OpenStudy (anonymous):

solve the equation. (If there is no solution, enter NO SOLUTION.) 3|x + 8| + 2 = 17 x = ___ (smaller value) x = ___ (larger value)

OpenStudy (anonymous):

subtract 2, divide by 3 to get \[|x+8|=5\] then solve \[x+8=5\] and \[x+8=-5\]

myininaya (myininaya):

3|x+8|+2-2=17-2 3|x+8|=15 |x+8|=5 so we have x+8=5 or x+8=-5 x=5-8 x=-5-8 x=-3 x=-13 -------------------------------- Check: x=-3 3|-3+8|+2=3*5+2=17 so x=-3 works! x=-13 3|-13+8|+2=3*5+2=17 so x=-13 works!

OpenStudy (anonymous):

thanks so much!

myininaya (myininaya):

np

OpenStudy (anonymous):

can you help with this one? Solve the equation. (If there is no solution, enter NO SOLUTION.) |x − 7| = |3x + 8|